@c -*-texinfo-*- @comment @c smallbook @setfilename calc.info @c [title] @settitle GNU Emacs Calc 2.02 Manual @setchapternewpage odd @comment

@tex @end tex

@c Fix some things to make math mode work properly. @iftex @textfont0=@tenrm @font@teni=cmmi10 scaled @magstephalf @textfont1=@teni @font@seveni=cmmi7 scaled @magstephalf @scriptfont1=@seveni @font@fivei=cmmi5 scaled @magstephalf @scriptscriptfont1=@fivei @font@tensy=cmsy10 scaled @magstephalf @textfont2=@tensy @font@sevensy=cmsy7 scaled @magstephalf @scriptfont2=@sevensy @font@fivesy=cmsy5 scaled @magstephalf @scriptscriptfont2=@fivesy @font@tenex=cmex10 scaled @magstephalf @textfont3=@tenex @scriptfont3=@tenex @scriptscriptfont3=@tenex @textfont7=@tentt @scriptfont7=@tentt @scriptscriptfont7=@tentt @end iftex

@c Fix some other things specifically for this manual. @iftex @finalout @mathcode‘@:=‘@: @c Make Calc fractions come out right in math mode @tocindent=.5pc @c Indent subsections in table of contents less @rightskip=0pt plus 2pt @c Favor short lines rather than overfull hboxes @tex @end tex @newdimen@kyvpos @kyvpos=0pt @newdimen@kyhpos @kyhpos=0pt @newcount@calcclubpenalty @calcclubpenalty=1000 @newcount@calcpageno @newtoks@calcoldeverypar @calcoldeverypar=@everypar @everypar=@calceverypar@the@calcoldeverypar @ifx@turnoffactive@undefinedzzz@def@turnoffactive@fi @ifx@ninett@undefinedzzz@font@ninett=cmtt9@fi @catcode‘@¯0